Two spherical masses each of mass 2kg are connected by a loght rod of lenght 50cm,calculate its rotational inertia.



(i) about an axis normal to it through iys center



(ii) about an axis at one of its ends

1
Expert's answer
2022-01-20T10:08:28-0500

1.

"I=\frac{md^2}2=0.25~kg\cdot m^2,"

2.

"I=\frac{2md^2}3=0.33~kg\cdot m^2."